OpenStudy (anonymous):

a. Find the surface area of the cube. b. Find the volume of the cube. the cube is 4x-5 wide.

OpenStudy (anonymous):

Let side of the cube 4x-5 = a So Surface Area = 6a^2 ==> 6(4x-5)^2 ==> 6(16x^2 - 40x + 25) ==> 96x^2 - 240x + 150 And Volume = a^3 ==> (4x - 5)^3 ==> 64x^3 - 240x^2 + 225x - 125